# Towards universal property prediction in Cartesian space: TACE is all you need

### v1

@@ 2.3.2 Element embedding
In the TACE framework, we denote node features by $\prescript{\nu_{1}}{c\,}{\mathbf{h}_{~i}^{(t)}}$, where $i$ refers to the atom index, $\nu_{1}\in\mathrm{L_{max}}$ indicates the tensor rank, $p$ indexes the contributing paths, $c$ denotes the feature channel, and $t$ represents the layer index. The rank-0 node features are initialized by applying a linear transformation to the one-hot encoding of the atomic species. Higher-rank node features are subsequently constructed via tensor contraction involving edge attributes. Node features will be iteratively updated through message passing mechanisms across successive layers. The prevailing MLIPs predominantly adopt one-hot encoding to scalar features for reasons analogous to those in large language models (LLMs). Elements are not entirely independent, and those within the same period or group often exhibit similar properties. This correlation allows the learned element embeddings to be linearly reduced in dimensionality and visualized after training, thereby revealing underlying relationships.
@@ 2.3.3 Universal invariant embedding
Here, we proposed the incorporation of additional graph- or node-level invariant physical quantities such as charges and total charge. For discrete element-like properties, we employ one-hot encoding followed by a linear layer, while for continuous properties we adopt an MLP-based embedding. All non-element embeddings are concatenated and then projected into the channel dimension. We denote this combined representation as universal invariant features, which is then directly added to the node features. This embedding approach allows us to incorporate any invariant quantity that can reasonably influence the (potential energy surface) PES. By embedding these properties, the model can account for their effects on the PES, improving accuracy and transferability across different chemical environments.
@@ 2.3.4 Universal equivariant embedding
Meanwile, we propose a general scheme for embedding equivariant physical quantities within the ACE framework. This approach is universally applicable to other models based on ACE, allowing any physically meaningful equivariant feature to be incorporated into the representation.
External field: The diverse responses of materials to external stimuli, ranging from linear to nonlinear and coupled effects, are crucial in determining the functional properties of a wide range of systems, including dielectrics, ferroelectrics, multiferroics, and piezoelectrics. Taking the external electric field as a representative example, we illustrate how such a field can be embedded into models based on ACE. Under universal equivalent embeddings, the $\mathbf{A}$-basis in Eqn. 10 is modulated by element- and channel-dependent weights $\mathrm{w_{c,z_{i}}}$, together with optional statistical information $\sigma$ about the field. This operation is applied after constructing the atomic basis and before forming the product basis. The atomic basis under an external electric field is then given by:
@@
Hereafter, the Field-Induced model is referred to as FITACE.
Magnetic moments: Similarly, the embedding scheme can be applied to both collinear and non-collinear magnetic systems. The overall procedure remains essentially the same, with the main distinction that the magnetic moments are treated independently for each atom. Hereafter, this magnetic model is referred to as mTACE.
@@ 2.3.5 Radial embedding
$j_{0}^{n}(r_{ij})=\sqrt{\frac{2}{c}}\frac{\sin(\frac{n\pi}{c}r_{ij})}{r_{ij}}f_{cut}(r_{ij})$ (6)
The edge length $r_{ij}=\|\mathbf{r}_{ij}\|$ is first expanded using a set of radial basis functions and subsequently multiplied by an envelope cutoff function. By tuning the hyperparameters of polynomial basis, we can effectively control the decay behavior of atomic contributions. Further details can be found in the work of ref. [30]. The resulting radial embeddings are subsequently passed through MLPs to generate path weights. We also implement $j_{n}$ of any order in the code.
@@ 2.3.6 Angular embedding
The angular information of an atomic environment are encoded in the normalized edge vector, $\hat{\mathbf{r_{ij}}}=\mathbf{r_{ij}}/r_{ij}$. In contrast to most of the SOAT models which rely on projections onto spherical harmonics , we adopt a cartesian representation. Specifically, we construct edge attributes $\prescript{(\nu_{2};\nu_{2};1)}{}{\mathbf{E}_{ij}}$, where $\nu_{2}\in\ell_{max}$ as:
@@ 2.3.7 Atomic basis
We begin by constructing the one particle basis $\prescript{\tilde{\nu_{3}}}{p,c}{\boldsymbol{\phi}}_{ij}^{(t)}$ where $\tilde{\nu_{3}}\in\mathrm{L_{max}}$ in Eqn. 9 via tensor contraction between node features and edge attributes. Each path is modulated by a learnable radial weight. The particle basis indeed captures pairwise ($2$-body) interactions at the first layer and higher-body interactions at the rest. The $\mathbf{A}$-basis in Eqn. 10 is then formed by aggregating the particle basis over neighboring atoms, followed by a decomposition to get highest-weight irreducible components. This representation is subsequently normalized by the average number of neighbors $\lvert{\mathcal{N}(i)}\rvert$} to ensure scale invariance. Our framework accommodates choice on distinct channels for atomic and product basis construction, while permitting the use of ICTs with $\ell<\tilde{\nu_{3}}$.
@@ 2.3.8 Product basis
Once the atomic basis is obtained, $\mathbf{B}$-basis in Eqn. 11 can be constructed via self-contraction. The n-correlation $\prescript{\nu_{3}}{}{\mathbf{B}_{i,n}}$ where $\nu_{3}\in\mathrm{L_{max}}$ is obtained via n times contracton with $\mathbf{A}$-basis itself. Note that after every tensor contracion, we could apply decomposition matrix to obtain highest weight.
@@ 2.3.9 Message passing
In TACE, the combination of equivariance, path and higher-body-order interactions allows the model to achieve the SOAT accuracy with only two message passing layers (T=2), outperforming Cartesian models such as CMAP [15] and HotPP [13] typically with T >=2>=2.
$\prescript{\nu_{3}}{c\,}{\mathbf{h}_{i}^{(t+1)}}=\begin{cases}\sum_{np_{2}c}W_{np_{2}cz_{i}\nu_{3}}^{(t)}\prescript{\nu_{3}}{p_{2}c\,}{\mathbf{B}}_{i,n}^{(t)}&\text{if }(\mathrm{t=0}~\&~\mathrm{T>1})\\ \sum_{np_{2}c}W_{np_{2}cz_{i}\nu_{3}}^{(t)}\prescript{\nu_{3}}{p_{2}c\,}{\mathbf{B}}_{i,n}^{(t)}+\sum_{\tilde{c}}W_{z_{i}\tilde{\nu_{3}}\tilde{c}c}^{(t)}\prescript{\tilde{\nu_{3}}}{c\,}{\mathbf{A}}_{i}^{(t)}&\text{if }(\mathrm{t>1}~\&~\mathrm{T>1})\text{ or }(\mathrm{T=1})\end{cases}$ (12)
The node features at layer $(t+1)$ are computed according to Eqn. 12. It is important to note that once the $\mathbf{A}$ and $\mathbf{B}$ basis are constructed, the subsequent feature update involves only learnable linear operations. As a result, the tensor ranks satisfy $\nu_{1}=\nu_{3}=\tilde{\nu_{3}}$. Moreover, the first message passing layer lacks self-connections except $\text{T}=1$, and thus does not incorporate residual links [33]. To enhance the expressive power of the model, we employ coupled features. The TACE framework further provides flexibility to choose between coupled features and element-dependent weights.
@@ 2.3.10 Readout
Both the dipole moment and polarizability are obtained using only linear operations. Specifically, the dipole moment requires only rank-1 tensors, while the polarizability is constructed from a combination of rank-0 and rank-2 tensors to yield a fully symmetric tensor. While, indeed, the inclusion of additional nonlinear gating operations, such as self-scaling based on the norm of the tensor itself or gating through scalar quantities, the model can, in principle, be extended to predict arbitrary-rank Cartesian tensors while preserving a body-ordered structure. This generalization enables the learning of complex tensorial properties, within a physically consistent framework. Finally, we emphasize that the energy prediction employs the isolated atomic energy either obtained as virtual quantities through least squares fitting or provided directly by the user from DFT calculations as the baseline. In addition, when universal invariant features are available, we can further obtain a baseline energy by passing them through a linear layer.
@@ 2.3.11 Charges equilibration
In principle, for any node-level rank-0 tensor that must be conserved, the surplus part of the prediction can be uniformly redistributed among atoms. As an illustrative example, we consider the case of atomic charges. While such uniform redistribution is possible, a more physically constrained choice is the Charge Equilibration (QEq) approach [34, 35]. To this end, we introduce two additional rank-0 readout heads to predict the electronegativity $\chi$ and hardness $\eta~\mathrm{(>0)}$ or $\frac{1}{\eta}$ for each atom. The atomic charges are then obtained by enforcing charge conservation through the following expression:
@@ 2.3.12 Latent Ewald Summation
where the complementary error function $\varphi(r)=\mathrm{erfc}(\frac{r}{\sqrt{2}\sigma})$. Hereafter, the model employing LES as a plugin is referred to as TACE-LES.
2.4 Lower weight and the TMP architecture
The primary model discussed in this work is TACE. However, as part of our unified Cartesian framework design, we highlight a key difference from spherical harmonics: the use of lower weights. The rationale behind employing lower weights is that they allow us to generate multiple irreducible components with only a single tensor contraction. In addition, within the Cartesian framework, we introduce another architecture, the Tensor Moment Potential (TMP). The most significant difference lies in the removal of the product basis. Instead, we employ layers T >> 2 to capture many-body information and add optional gating operations. This alternative is slightly less accurate, but generally at least twice as fast, since in Cartesian space, under our current implementation, self-contraction operations are slower than spherical harmonics. Moreover, TMP also avoids several issues that are inherent to the ACE formalism. As it likewise constructs atomic basis functions, all functionalities described earlier are retained. More detailed evaluations and improvements will be left for future work.
In the current SOAT model, each tensor-product path is associated with a radial weight. Due to the presence of multiple channels, each edge carries rich radial information. The previously mentioned TACE model is a typical example of such embedding. However, in the Cartesian space, through ICTD, for generic Cartesian tensors of rank 0 to 4, the number of weights is 1, 1, 3, 7 and 19 respectively. This indicates that, compared to spherical harmonics, Cartesian tensors not only provide computational advantages for rank $\leq 4$, but also, from an intuitive standpoint, using low-weight tensors should not result in any loss of accuracy.
We argue that the most effective way to embed low weights is to transform low-weight tensors into components with the same rank and weight within the convolution. This is straightforward for tensors of rank $\leq 2$, but for rank $\geq 3$ tensors, a direct approach does not provide a simple conversion. Therefore, in the following discussion, our goal is to demonstrate the independence of low-weight tensors as irreducible representations, as well as their potential applications.
We employ ICTD in two ways: the first is “max”, which uses only the highest weight for a given rank; the second is “all”, which utilizes all available weights. For the “max” approach, we do not elaborate further here. In the “all” case, we do not directly embed radial weights along the path; however, each path is still modulated by a learnable weight initialized to 1.0. Instead, after tensor contraction, the messages obtained on each edge are processed via ICTD, at which point radial information is embedded through $\sum_{\ell,q}R_{\ell,q}\cdot\prescript{{(\nu;\ell;q)}}{}{\mathbf{T}}$. In other words, the number of radial weights depends on the tensor rank rather than the possibile path. The “all” strategy does not necessarily yield better performance, as it is not our optimal implementation. Nonetheless, pairing a weight $\ell$ with a learnable radial weight is a natural and straightforward approach. Finally, we emphasize that our tensor contraction operations do not require the input tensors to be irreducible, reducible cases can be handled in a unified manner.
2.5 Multi-fidelity training — Urea
From this section onward, we will present benchmark results for various systems, starting with some special benchmarks that we consider particularly important, such as multi-fidelity training, spectra, Hessians, magnetic systems, and external field responses, as well as charge-related properties. We will then move on to more conventional benchmarks, such as in-domain rMD17 and out-of-domain 3BPA. Finally, we will discuss the advantages and disadvantages of ACE, and present the application results of TACE in heterogeneous catalysis. We consider the urea molecule from the QeMFi [36] dataset to demonstrate the effect of universal embedding on discrete variables. Five levels of theory are employed: STO-3G, 3-21G, 6-31G, def2-SVP, and def2-TZVP, each with 15,000 configurations. STO-3G is used as the baseline with a train:valid:test split of 41:9:50, while the other levels adopt a split of 1:9:90. In total, six models are trained: five using datasets from each level individually and one using a combined dataset incorporating all levels. Comparisons are made on the test set. As shown in Fig. 1, the multi-fidelity model achieves excellent accuracy even when only 2.2% of the training data are from high-fidelity levels.
The trends in the columns of Fig. 1 are consistent with expectations. Joint training across multiple basis sets may slightly degrade the accuracy at lower levels, but the improvement at higher levels is substantial. Since this example might be somewhat extreme, we also examined cases with an increased proportion of high-fidelity data, with the corresponding results provided in the SI. In these cases, the error at high-fidelity levels decreases rapidly as the size of the training set increases. Nevertheless, the root-mean-square error of the multi-fidelity model is approximately half that of a single-level model.
Figure 1: Comparison of Multi-fidelity and Single-fidelity Training. a-e Test set results of models trained using multi-fidelity (mixed-basis) data. f-j Results of models trained separately with each individual basis set. Panels in the same column correspond to the same test set.
2.6 Spectra — Water
As an illustrative application to complex molecular liquids, we assess the performance of TACE on a dataset comprising 1593 configurations of liquid water [37], computed at the revPBE0-D3 level, which is known to provide a reliable description of the structure and dynamics of water across a range of pressures and temperatures [38].
The RMSE of energy per atom and forces are listed in Table 1. Invariant models, such as BPNN and DeepMD [2, 6], rely on lower-body descriptors without message passing yield comparable errors, which rank among the highest in this comparison. In contrast, the linear ACE incorporates higher-order features without message passing are better. The best-performing models include those that combine higher-body-order representations with equivariant message passing, such as NequIP, MACE and CAMP [7, 9, 15], all of which achieve significantly lower errors, demonstrating the benefits of incorporating both equivariance and complex many-body correlations. Our TACE model delivers the highest accuracy on both energy and forces predictions. Interestingly, similar to ref. [39], a small-scale TACE model with invariant messages ($\mathrm{L_{max}=0}$) already achieves lower errors than other top-performing models, and increasing the model size offers only modest gains in accuracy. To assess the stability of molecular dynamics simulations, we performed three independent 300-ps simulations of water at 300 K, 500 K, 1000 K, 1500 K, and 2000 K using the $\mathrm{L_{max}=0}$ model. The system simulated consists of 512 water molecules, corresponding to a density of 1 $\mathrm{g/cm^{3}}$. A time step of 1 fs was employed, and temperature control is maintained using a Nosé-Hoover thermostat. The O-O radial distribution function (RDF) is shown in Fig. 2 (a). At 300 K, the calculated RDF agrees remarkably well with experimental results from X-ray diffraction studies [40]. We also computed the mean square displacement (MSD) and obtained the diffusion coefficient using the Einstein relation, yielding $0.26(1),1.58(1),3.17(7),4.66(7),5.90(28)\times 10^{-5}\ \mathrm{cm^{2}/s}$ at 300, 500, 1000, 1500, 2000K, which is in excellent consistency with the AIMD value of $2.67\times 10^{-5}\ \mathrm{cm^{2}/s}$ in 300K [41]. The TACE potential maintains stable MD simulations up to 2000 K, highlighting its robustness. This stability enables both reliable statistical sampling and potential refinement via active learning, where configurations from MD are iteratively used for retraining.
The response of a material to mechanical, magnetic, or electric perturbations is characterized by tensorial response coefficients that inherently describe the directional dependence and coupling of physical properties. As an illustrative example, we predict the dipole moment and polarizability for various water systems. For the dipole moment, we directly predict a vector $\nu=1$. For the polarizability, since it is a traceless and fully symmetric tensor, we predict two components: the trace part ($\ell=0$) and the traceless fully symmetric part ($\ell=2$). The prediction accuracy is reported in terms of the relative RMSE [42] in Table 2, defined as the RMSE normalized by the intrinsic standard deviation of the validation set. The explicit formula can be found in [43]. We found that our model achieves very high accuracy in predicting dipole moments, reaching the best performance in three out of the four systems considered. Its performance in predicting polarizabilities is also outstanding. However, some anomalous values can be observed in the Table 2. For the invariant REANN model, dipole moments are predicted with satisfactory accuracy in finite systems. However, in extended systems, the approach of deriving dipole moments from charges and center of mass may present certain limitations. In the case of TACE, we observe relatively larger errors in polarizability prediction for liquid water compared with other advanced models, which is unexpected given its strong performance across other systems. Closer inspection reveals that the polarizability labels are not strictly fully symmetric tensors but are instead indirectly obtained through the tensorial extension of the Clausius-Mossotti relation [44] from the instantaneous dielectric response tensor of liquid water configurations. We find that HotPP implementation treats polarizability as a reducible Cartesian tensor, predicting the $\ell=0$ part alongside a general rank-2 tensor, and subsequently applying manual symmetrization. By contrast, our model enforces full tensor symmetry by explicitly summing the $\ell=0$ and $\ell=2$ components. From a theoretical standpoint, we consider our formulation to be more appropriate and thus accordingly attribute the comparatively poor polarizability performance in liquid water to limitations in the intrinsic noise of dataset labels [42]. To obtain the infrared and Raman spectra, we employed an NPT-equilibrated configuration of 512 water molecules from ref. [13]. This was followed by an additional 100 ps equilibration in the NVT ensemble using TACE trained on the dataset of ref. [45]. Subsequently, three production simulations were carried out in the NVE ensemble, each with a duration of 200ps. Throughout all simulations, a time step of 0.5 fs was applied. The infrared and Raman spectra were then obtained by performing Fourier transforms of the respective autocorrelation functions. It should be emphasized that quantum effects in water can significantly influence both the RDF and the simulated spectra, as demonstrated in refs. [46, 47, 48, 38].

### v2

@@ 2.3.2 Element embedding
In the TACE framework, we denote node features by $\prescript{\nu_{1}}{c\,}{\mathbf{h}_{~i}^{(t)}}$, where $i$ refers to the atom index, $\nu_{1}\in\mathrm{L_{max}}$ indicates the tensor rank, $p$ indexes the contributing paths, $c$ denotes the feature channel, and $t$ represents the layer index. The rank-0 node features are initialized by applying a linear transformation to the one-hot encoding of the atomic species. Higher-rank node features are subsequently constructed via tensor contraction involving edge attributes. Node features will be iteratively updated through convolution across successive layers. The prevailing MLIPs predominantly adopt one-hot encoding to scalar features for reasons analogous to those in large language models (LLMs). Elements are not entirely independent, and those within the same period or group often exhibit similar properties. This correlation allows the learned element embeddings to be linearly reduced in dimensionality and visualized after training, thereby revealing underlying relationships.
@@ 2.3.3 Universal invariant embedding
Here, we proposed the incorporation of additional graph- or node-level invariant physical quantities such as charges and computational levels. For discrete element-like properties, we employ one-hot encoding followed by a linear layer, while for continuous properties we adopt an MLP-based embedding. All non-element invariant embeddings are concatenated and then projected into the channel dimension. We denote this combined representation as universal invariant features, which is then directly added to the node features. This embedding approach allows us to incorporate any invariant quantity that can reasonably influence the PES. By embedding these properties, the model gains additional descriptive flexibility for the PES, thereby improving its accuracy and transferability across different chemical environments.
@@ 2.3.4 Universal equivariant embedding
Meanwhile, we propose a general scheme for embedding equivariant physical quantities within the ACE framework. This approach is universally applicable to other ACE-based models, allowing any physically meaningful equivariant feature to be incorporated into the representation. Strictly speaking, pseudovectors need to be considered within the O(3) framework and should be embedded in the corresponding irreducible components. Additionally, if time-reversal symmetry is taken into account, our treatment needs to be modified accordingly; however, from a numerical perspective, fitting remains feasible.
External electric field: The diverse responses of materials to external stimuli, ranging from linear to nonlinear and coupled effects, are crucial in determining the functional properties of a wide range of systems, including dielectrics, ferroelectrics, multiferroics, and piezoelectrics. Taking the external electric field as a representative example, we illustrate how such a field can be embedded into models based on ACE. Under universal equivariant embeddings, the $\mathbf{A}$-basis in Eqn. 11 is modulated by element- and channel-dependent weights $\mathrm{w_{c,z_{i}}}$, together with optional statistical information $\sigma$ about the field. This operation is applied after constructing the atomic basis and before forming the product basis. The atomic basis under an external electric field is then given by:
@@
Magnetic moments: Similarly, the embedding scheme can be applied to both collinear and non-collinear magnetism. It is important to emphasize that the physical symmetries associated with spin degrees of freedom differ between the two cases. In non-collinear cases, geometric tensors transform according to SO(3) under spatial rotations, while spinors transform according to SU(2), with these transformations intricately coupled. In contrast, collinear cases define a global spin axis, rendering the transformations of atomic positions and spins effectively separable. Therefore, in principle, spins in collinear setups should be treated as scalars. For non-collinear setups, the overall procedure remains essentially the same as in the presence of an external field, with the main distinction that the magnetic moments are treated for each atom.
@@ 2.3.5 Radial embedding
$j_{0}^{n}(r_{ij})=\sqrt{\frac{2}{c}}\frac{\sin(\frac{n\pi}{c}r_{ij})}{r_{ij}}f_{\mathrm{cut}}(r_{ij})$ (7)
The edge length $r_{ij}=\|\mathbf{r}_{ij}\|$ is first expanded using a set of radial basis functions and subsequently multiplied by an envelope cutoff function to ensure the smoothness of the PES. By tuning the hyperparameters of polynomial basis, we can effectively control the decay behavior of atomic contributions. Further details can be found in the work of ref. [33]. The resulting radial embeddings are subsequently passed through MLPs to generate path weights. We also implement $j_{n}$ of any order in the code. In our trials, using the first few orders did not result in a significant improvement or degradation of performance, though it may prove useful in future applications.
@@ 2.3.6 Angular embedding
The angular information of an atomic environment are encoded in the normalized edge vector, $\hat{\mathbf{r_{ij}}}=\mathbf{r_{ij}}/r_{ij}$. In contrast to most of the SOTA models which rely on projections onto spherical harmonics , we adopt a cartesian representation. Specifically, we construct edge attributes $\prescript{(\nu_{2};\nu_{2};1)}{}{\mathbf{E}_{ij}}$, where $\nu_{2}\in\ell_{max}$ as:
@@ 2.3.7 Atomic basis
We begin by constructing the one particle basis $\prescript{\tilde{\nu_{3}}}{p,c}{\boldsymbol{\phi}}_{ij}^{(t)}$ where $\tilde{\nu_{3}}\in\mathrm{L_{max}}$ in Eqn. 10 via tensor contraction between node features and edge attributes. Each path is modulated by a learnable radial weight. The particle basis indeed captures pairwise ($2$-body) interactions at the first layer and higher-body interactions at the rest. The $\mathbf{A}$-basis in Eqn. 11 is then formed by aggregating the particle basis over neighboring atoms, followed by an optional decomposition to get highest-weight irreducible components. It should be noted that the normalization constant $\mathcal{N}(i)$ can be dynamically determined based on node features, node attributes or edge features, or it can be set according to the average number of neighboring atoms. Our framework also accommodates choice on distinct channels for atomic and product basis construction, while permitting the use of ICTs with $\ell<\tilde{\nu_{3}}$.
@@ 2.3.8 Product basis
Once the atomic basis is obtained, $\mathbf{B}$-basis in Eqn. 12 can be constructed via self-contraction. The n-correlation $\prescript{\nu_{3}}{}{\mathbf{B}_{i,n}}$ where $\nu_{3}\in\mathrm{L_{max}}$ is obtained via n times contracton with $\mathbf{A}$-basis itself. Note that after every tensor contracion, we could apply an optional decomposition matrix to obtain highest weight.
@@ 2.3.9 Message passing
In TACE, the combination of equivariance, path and higher-body-order interactions allows the model to achieve the SOTA accuracy with only two layers (one message passing layers and T=2), outperforming Cartesian models such as CMAP [17] and HotPP [15] typically with T >2>2.
$\prescript{\nu_{3}}{c\,}{\mathbf{h}_{i}^{(t+1)}}=\begin{cases}\sum_{np_{2}c}W_{np_{2}cz_{i}\nu_{3}}^{(t)}\prescript{\nu_{3}}{p_{2}c\,}{\mathbf{B}}_{i,n}^{(t)}&\text{if }(\mathrm{t=0})\text{ and }(\mathrm{T>1})\\ \sum_{np_{2}c}W_{np_{2}cz_{i}\nu_{3}}^{(t)}\prescript{\nu_{3}}{p_{2}c\,}{\mathbf{B}}_{i,n}^{(t)}+\sum_{\tilde{c}}W_{z_{i}\tilde{\nu_{3}}\tilde{c}c}^{(t)}\prescript{\tilde{\nu_{3}}}{c\,}{\mathbf{A}}_{i}^{(t)}&\text{if }(\mathrm{t>0})~~~\text{or }(\mathrm{T=1})\end{cases}$ (13)
The node features at layer $(t+1)$ are computed according to Eqn. 13. It is important to note that once the $\mathbf{A}$ and $\mathbf{B}$ basis are constructed, the subsequent feature update involves only learnable linear operations. As a result, the tensor ranks satisfy $\nu_{1}=\nu_{3}=\tilde{\nu_{3}}$. Moreover, the first message passing layer lacks self-connections except $\text{T}=1$, and thus does not incorporate residual links [36]. To enhance the expressive power of the model, we employ coupled features. The TACE framework further provides flexibility to choose between coupled features and element-dependent weights.
@@ 2.3.10 Readout
For tensors that are directly predicted, such as the dipole moment, polarizability, direct forces, direct stress, and so on, we only need to add a readout corresponding to the appropriate rank. Taking the polarizability as an example, it is constructed from a combination of rank-0 and rank-2 tensors to yield a fully symmetric tensor. Alternatively, in the reducible case, it can be obtained directly using ICTD. Finally, we emphasize that the energy prediction employs the isolated atomic energy either obtained as virtual quantities through least squares fitting or provided directly by the user from DFT calculations as the baseline. In addition, when universal invariant features are available, we can further obtain an optional baseline energy by passing them through a linear layer.
@@ 2.3.11 Charges equilibration
In principle, for any node-level rank-0 tensor that must be conserved, the surplus part of the prediction can be uniformly redistributed among atoms. As an illustrative example, we consider the case of atomic charges. While such uniform redistribution is possible, a more physically constrained choice is the Charge Equilibration (Qeq) approach [37, 38]. To this end, we introduce two additional rank-0 readout MLP to predict the electronegativity $\chi$ and hardness $\eta~\mathrm{(>0)}$ or $\frac{1}{\eta}$ for each atom. The atomic charges are then obtained by enforcing charge conservation through the following expression:
@@ 2.3.12 Latent Ewald Summation
where the complementary error function $\varphi(r)=\mathrm{erfc}(\frac{r}{\sqrt{2}\sigma})$. It should be noted that, depending on the functional form, LES can simultaneously describe both electrostatic and dispersion interactions. Hereafter, the model employing LES as a plugin is referred to as TACE-LES.
Models based on tensors have become widely adopted and extensively studied in recent years, providing a solid foundation for many applications in physics, chemistry, and biology. In contrast, Cartesian tensors, despite their natural and often more straightforward formulation in certain contexts, have not yet been presented in a systematic and comprehensive manner in the existing literature. To address this gap, we begin by reviewing the fundamental concepts of Cartesian tensors (2.2 and 2.3), which serve as the theoretical basis for our approach. We then proceed to describe our method in detail, highlighting its distinguishing features and advantages over existing frameworks.
Table 1: RMSE for energy (E, meV) and forces (F, meV/Å) for the 3BPA dataset. The best-performing results are both bolded and underlined.00footnotetext: All models were trained using dataset at $\mathrm{T}=300\,\mathrm{K}$. Results for all models except TACE are taken from ref. [16]. Here, both TACE models achieve excellent accuracy. However, note that the use of “all” is merely intended to demonstrate its potential applicability, and does not represent the optimal way of employing it. We only show the performance of “all” here, while all other benchmarks are based on “max”.
GAP[39] ACE[23] CACE[16] Allegro[9] NequIP[8] MACE[10] $\mathbf{TACE_{all}}$ $\mathbf{TACE_{max}}$
300K E 22.8 7.1 6.3 3.8 3.3 3.0 2.87 2.69
F 87.3 27.1 21.4 13.0 10.8 8.8 9.84 8.68
600K E 61.4 24.0 18.0 12.1 11.2 9.7 9.20 8.10
F 151.9 64.3 45.2 29.2 26.4 21.8 22.59 19.88
1200K E 166.8 85.3 58.0 42.6 38.5 29.8 31.90 23.42
F 305.5 187.0 139.6 83.0 76.2 62.0 63.10 54.26
dihedral E - - - - 23.2 7.8 13.82 11.00
F - - - - 23.1 16.5 18.99 16.45
2.4 Out-of-Domain — 3BPA
[…]
Across all splits, TACE attains force and energy RMSE competitive with or better than leading spherical-harmonic equivariant models (NequIP, Allegro, MACE), while using only 64 channels compared to 256 in a comparable MACE baseline (Table 1.) Moreover, the 3BPA dataset is employed in our study to investigate the differences between coupled and uncoupled feature representations in Eqn. 13. We observe that coupled features consistently achieve higher accuracy on force predictions across all test sets. Nonetheless, this result indicates that Cartesian models may achieve higher accuracy with fewer channels when using coupled channel features and if lower-weights can be effectively utilized, the channel can be further reduced.
Figure 2: Comparison of Different Mixed Precision Training Approaches. From left to right, the basis sets are STO-3G, 3-21G, 6-31G, def2-SVP, and def2-TZVP. The test set is the same for each column. a-e show TACE using single-fidelity training. f-j show TACE using multi-fidelity training. k-o show TACE using multi-head training, p-t show MACE using multi-head training. u-y show TACE using both multi-fidelity and multi-head schemes.
2.5 Mixed Precision Training — Urea
Training on datasets with mixed precision has emerged as an important strategy to reduce the cost of generating high-accuracy reference data. Here, we treat the computational level as an additional discrete label using the universal invariant embedding framework and train TACE on datasets of different electronic-structure fidelities within a single unified model. This strategy ,which we refer as multi-fidelity training, differs conceptually from standard multi-head training, implemented in MACE [10, 40]. In multi-head training, separate output heads are attached for each fidelity level while sharing a common representation; in multi-fidelity training, the fidelity label is embodied directly into the representation itself, so that the internal features become level-aware. Both strategies can coexist, but multi-fidelity training offers a more natural route to cross-fidelity generalization.
We evaluate this idea on the QeMFi [41] dataset to compare the effects of single-fidelity (SF), multi-fidelity (MF), multi-head (MH), and their combined approaches on predictive performance. The dataset contains 15,000 molecular configurations for each of five levels, computed using TD-DFT with the CAM-B3LYP functional. The basis sets, in order of increasing accuracy, are: STO-3G, 3-21G, 6-31G, def2-SVP, and def2-TZVP. A large low-fidelity dataset supplemented with a small amount of high-fidelity data: STO-3G serves as the baseline with a train:validation:test split of 41:9:50, while the other levels use a split of 1:9:90, so that only a very small amount of the total training data comes from the higher levels. Single-fidelity models are trained using datasets from only one basis set, whereas other models are trained on a combined dataset incorporating all five fidelity levels. As shown in Fig. 2, the red region indicates that, for the high-fidelity dataset, the model trained solely on limited high-fidelity data fails to generalize to unseen configurations. In contrast, both multi-fidelity and multi-head training greatly reduce these errors as high-fidelity supervision is now supported by abundant low-fidelity structures. Multi-fidelity TACE achieves the lowest errors across high-energy configurations, outperforming both the multi-head variant of TACE and the multi-head MACE baseline under matched architectures and hyperparameters. This improvement mainly arises from use of coupled features, which improves accuracy under the same channel size. Although the yellow region in Fig. 2 does not show a significant advantage from combining multi-head and multi-fidelity schemes, this combination still provides a more flexible modeling option. It is also worth noting that the introduction of high-fidelity data may slightly degrade the accuracy on low-fidelity datasets, but the gain at the high-fidelity levels is substantial. Overall, these experiments demonstrate that universal embeddings let TACE act as a cross-fidelity learner.